OpenStudy (anonymous):

Marie has $830 in her bank account and withdraws $60 each month. Denise has $970 in her bank account and withdraws $80 each month. In how many months will Marie and Denise have equal amounts of money in their accounts? A. 4 B. 5 C. 7 D. 9

Directrix (directrix):

Marie has 830 and removes 60 per month. 830 - 60*m is the amount Marie has after m months Denise has 970 and removes 80 per month. 970 - 80*m is the amount Denise has after m months.

Directrix (directrix):

@raventayler13 Solve this equation for m. 830 - 60*m = 970 - 80*m

Directrix (directrix):

Post what you get, okay?

OpenStudy (anonymous):

ok i think i got it is this it? Simplifying 830 + -60m = 970 + -80m Solving 830 + -60m = 970 + -80m Solving for variable 'm'. Move all terms containing m to the left, all other terms to the right. Add '80m' to each side of the equation. 830 + -60m + 80m = 970 + -80m + 80m Combine like terms: -60m + 80m = 20m 830 + 20m = 970 + -80m + 80m Combine like terms: -80m + 80m = 0 830 + 20m = 970 + 0 830 + 20m = 970 Add '-830' to each side of the equation. 830 + -830 + 20m = 970 + -830 Combine like terms: 830 + -830 = 0 0 + 20m = 970 + -830 20m = 970 + -830 Combine like terms: 970 + -830 = 140 20m = 140 Divide each side by '20'. m = 7

Directrix (directrix):

Yes. Which website is that from? There is mathpapa.com and one with a pirate. I'm just wondering.

OpenStudy (anonymous):

http://www.geteasysolution.com i like this website for these things when there's an ='s involved with something on the both sides i go to this site. @Directrix

Directrix (directrix):

Thanks. The site I used did not allow for copying and pasting. So, I've been looking for something better.

OpenStudy (anonymous):

oh well this is better but not all math problems can be solved on this site so i use 2 sites that one and https://mathway.com/ to solve problems. @Directrix

Directrix (directrix):

I'll check it, too. Thanks.

OpenStudy (anonymous):

welcome